### v3

@@ IV-A Overview
As specified in this work, we utilize the proposed CAMO-FS dataset containing images of camouflaged animals in the wild to establish the evaluation of our baseline and proposed improvement. We follow the concept procedure published in FSOD [37, 36, 38]. In the first stage of the base phase, we train our model with abundant data from 80 classes of the COCO dataset as proposed in [36]. In the second stage of the novel phase, we evaluate the performance of having $K=1,2,3,5$ shots per each novel class.
@@
Our MTFA [70] baseline is implemented using Detectron2 framework [73]. Our backbone is ResNet-101 [74] with Feature Pyramid Network [75]. Each experiment is set up with a single GPU GeForce RTX 2080Ti with a batch size of 2 images. The novel phase has a learning rate of 0.00125 inferred from the MTFA configuration. We set the balance parameters $\alpha$ = $1\times 10^{-1}$ and $\beta$ = $1\times 10^{-2}$ when we train the model with instance triplet and instance memory loss function, respectively. Please visit the publication [37] or [73] for more details on other parameters.
@@ IV-B Results and Discussion
State-of-the-art comparison. To prove the effectiveness of our proposed methods, we conducted experiments on our proposed CAMO-FS dataset. We tested with $K=\{1,2,3,5\}$ shots, respectively. Since several recent work have not published their source code [64, 63], we adopted the typical models addressing both detection and instance segmentation tasks to compare with our proposed methods. Table IV presents the evaluation of the performance of our methods of instance triplet loss and memory storage over our baseline MTFA [70], the model of Mask R-CNN [40] with sigmoid classifier, and the state-of-the-art method iFS-RCNN [61] in the approach of few-shot instance segmentation. We reported experiments on those models and chose the common COCO-80 ResNet-101 as their base model to apply our proposed methods. The details of this decision are declared in the ablation section. In terms of instance segmentation, we improved over MTFA [70], Mask ${}^{\dagger}$ [40], and iFS-RCNN [61] by getting average AP values of $6.23\%,8.04\%,7.13\%$, respectively thanks to instance triplet loss, and $7.35\%,7.96\%,6.25\%$, respectively thanks to instance memory storage. Regarding object detection, our FS-CDIS got average amounts of $7.14\%,7.34\%,6.96\%$, respectively with instance triplet loss and $7.34\%,7.45\%,5.88\%$, respectively with instance memory storage. The detailed performance of our methods is in Table IV. Despite the limited results, we defeated the very early models on detection and instance segmentation tasks on camouflaged images.
@@
Proposed modules evaluation. In Table V, we also present the results of the baseline MTFA [70] with its original default configuration along with our proposed improvements. On top of the baseline MTFA [70], we establish fine-tuning configuration on this model by training all heads of classification, box regression, and mask prediction on few-shot novel data. The reported results prove the performance of the proposed instance triplet loss and instance memory storage.
TABLE V: The improvement of our proposed instance triplet loss and instance memory storage over MTFA [37]. The best performance is marked in boldface. # denotes the Number of shots, “Memory” is Instance Memory Storage and “Triplet” is Instance Triplet Loss.
# Method AP AP50 AP75 APs APm APl AR1 AR10 ARs ARm ARl Instance Segmentation 1 Baseline MTFA 3.66 5.37 4.09 22.42 4.35 2.01 11.30 13.58 25.97 12.96 12.53 MTFA + Triplet 4.46 8.21 4.60 21.33 4.13 4.01 12.36 15.04 23.17 9.49 16.67 MTFA + Memory 5.46 9.20 6.17 27.79 6.20 4.01 17.08 19.99 29.41 11.45 20.89 2 Baseline MTFA 6.21 8.92 7.28 32.64 7.75 3.50 18.88 21.12 35.82 15.49 20.14 MTFA + Triplet 5.57 9.45 6.04 25.83 3.01 5.37 15.67 17.33 26.13 7.37 17.50 MTFA + Memory 6.95 10.72 7.60 33.62 5.73 6.44 20.00 22.15 34.25 13.86 20.92 3 Baseline MTFA 6.16 8.95 6.68 33.74 6.19 5.08 20.25 22.95 36.83 16.31 21.63 MTFA + Triplet 6.41 10.67 6.72 30.39 5.17 5.30 20.69 22.98 31.90 15.69 22.53 MTFA + Memory 7.36 11.23 8.49 37.03 6.24 5.64 24.40 27.69 38.44 17.02 26.71 5 Baseline MTFA 5.95 8.67 6.94 34.71 6.25 4.85 21.29 24.42 36.86 14.51 24.83 MTFA + Triplet 8.48 13.43 9.80 36.66 5.75 8.04 23.83 26.66 37.03 11.62 25.91 MTFA + Memory 9.61 14.61 11.73 38.60 5.79 10.40 26.65 30.37 39.21 12.26 30.02 Object Detection 1 Baseline MTFA 2.93 5.86 2.20 20.95 4.18 2.03 9.25 10.84 21.74 11.49 8.77 MTFA + Triplet 4.04 8.65 2.98 20.50 4.90 4.22 12.89 15.53 20.73 11.45 17.46 MTFA + Memory 4.50 9.14 3.45 22.88 5.61 3.54 13.14 15.22 23.14 8.78 16.33 2 Baseline MTFA 5.90 8.87 6.83 33.04 9.74 3.10 17.26 19.25 34.04 15.74 19.61 MTFA + Triplet 7.28 11.22 8.25 32.31 10.72 6.83 20.52 22.69 32.34 14.88 23.52 MTFA + Memory 6.95 10.88 7.75 33.93 7.49 6.81 19.84 22.01 34.10 15.04 21.47 3 Baseline MTFA 5.84 8.98 6.29 34.56 7.78 4.31 19.13 21.83 35.80 15.93 21.09 MTFA + Triplet 7.49 11.51 8.23 38.45 8.61 6.38 24.88 27.52 38.55 17.66 27.44 MTFA + Memory 7.55 11.45 8.50 38.07 9.21 5.70 24.20 27.29 38.50 18.10 27.56 5 Baseline MTFA 5.84 9.13 6.04 35.44 8.17 4.22 19.67 22.96 35.94 14.16 22.58 MTFA + Triplet 9.76 14.37 11.12 40.05 8.82 9.89 25.93 29.28 40.05 12.53 30.32 MTFA + Memory 10.36 16.27 11.79 39.32 8.08 11.36 26.34 30.30 39.35 12.37 30.91
@@
TABLE VI: Ablation study on the base model with 1-shot results. The best, and second best performances are marked in red, and blue, respectively. “Memory” is Instance Memory Storage and “Triplet” is Instance Triplet Loss.
@@
In general, our approaches achieve outstanding results in comparison with the baseline. Our improvements surpass MTFA by a remarkable margin. Regarding the instance segmentation, our method improves 1.9%, 3.5%, and 2.3% in terms of AP, AP@50, and AP@75, respectively. These results manifest the efficiency of our methods in the context of few-shot camouflaged instances. Both loss functions enhance the discrimination between foreground and background features which strongly supports the model to segment pixels that belong to the camouflaged animals. Regarding the memory bank and the triplet loss function, the results of the memory loss function are higher than those of the triplet loss function by about 1%. We realize that storing representatives for each class is a crucial element in few-shot learning. This technique not only expands the variants during training but also increases the consistency per class, thereby model can segment difficult objects better. In these ways, we also improve the corresponding results in camouflage object detection.
@@
In Table V, our improvements help the model segment animals in various sizes. Specifically, all three metrics including APs, APm, and APl improve in comparison with the baseline model, which demonstrates that our model well segments small, medium, and large animals. This situation also happens in the detection problem. When data is very scarce as in a 1-shot or 2-shot setting, the triplet loss function has comparative results with the memory loss function. However, in the context of 3-shot or 5-shot settings, the memory loss function demonstrates outstanding efficiency thanking to storing and updating the memory via iterations to create discriminative features on a global level. Figure 7 illustrates the qualitative comparison among the results of 5-shot settings of the baseline MTFA [70] and our proposed methods of Instance Triplet Loss and Instance Memory Storage. We chose to visualize the images with the confidence threshold of $0.5$, which released a huge number of predictions with low confidence from the models. The four final rows indicate exemplary cases that either FS-CDIS-ITL or FS-CDIS-IMS can figure out camouflaged instances compared to the baseline.
@@
TABLE VII: Ablation study on the margin and the $\alpha$ ratio of the instance triplet loss in 1-shot settings. The best, and second best performances are marked in red, and blue, respectively.
@@
Base model ablation study. We also conduct ablation experiments on different backbone base models of the COCO settings including general and few-shot concepts. To be detailed, we report the performance of our proposed method of instance triplet loss and instance memory storage over four different backbones. The considered backbones are ResNet-50 and ResNet-101 [74]. The two base datasets are MS-COCO with 80 classes and 60 classes, respectively. Thus, it led to the combination of four different base models (i.e. COCO-80/60 R-101, COCO-80/60 R-50). As can be seen from Table VI, the performance of applying COCO-80 R-101 base weight yields better results among others evaluated on AP, AP@50, and AP@75 in both segmentation and detection tasks. In both cases of our two proposed improvements, the ablation results demonstrate our selection of COCO-80 R-101 is the best among the tested backbones of the base phase. For the segmentation task, we achieve $4.46$ and $5.46$ of AP reported for triplet loss and memory storage, respectively. For the detection task, we reach $4.04$ and $4.50$ also of AP, respectively. In summary, the chosen backbone of the base weight presents a higher performance of around $1$% to $2$% of evaluated on common metrics as in the table. To be explained, the base from COCO-80 contains more semantic concepts in comparison with the COCO-60 base, which leads to the higher performance reported. Note that all the results in this ablation section are reported for the 1-shot setting.
Ablation on instance triplet loss component. In terms of the instance triplet loss described in Eq. 1, we establish ablation experiments to evaluate the performance of the model with different configurations of margin and $\alpha$ value. To this end, we set up the margin varying from 0 to 1, with a step of 0.25. For the $\alpha$ ratio of the loss function (Eq. 3), we check out $\alpha$ = {$1,1\times 10^{-1},1\times 10^{-2}$}. To be enhanced, the margin value indicates how distinguished foreground and background features are. Meanwhile, the $\alpha$ controls the effect of the instance triplet loss on the total loss of the framework. Table VII presents the evaluation of both detection and segmentation issues in 1-shot manner. As can be inferred from the table, the effect of $\alpha$ decides which margin should be selected for the triplet loss. With $\alpha=1$ meaning we keep the original ratio of the loss, the segmentation result in 1-shot setting yields the highest performance of 5.16% mAP with a 0.75 margin value. Meanwhile, the detection result gets the highest performance of 4.36% with $\alpha=1\times 10^{-1}$ and zero margin. This table offers a better understanding of the impact of $\alpha$ and the margin over the total performance.
TABLE VIII: Ablation study on the capacity of the instance memory storage in 1-shot settings. The best, and second best performances are marked in red, and blue, respectively.
@@
Ablation on instance memory storage component. As for the instance memory storage, as introduced in Eq. 2, and Eq. 3, there are several parameters that need analyzing, listed as the amount of capacity in the memory storage and the $\beta$ ratio controlling the effect of the memory storage loss in the total loss. Table VIII, and Table IX present the ablation experimental results of those issues, respectively. In terms of the capacity of the memory storage, we establish experiments on a range of memory capacity of $2^{i}$ where $i=\{5,6,7,8,9,10\}$. The reported results figure out that the performance on both segmentation and detection tasks increases with a larger capacity of memory storage. To be detailed, with a capacity of 512, the mAP metric achieves the highest value among configurations, i.e. $4.76\%$ and $4.48\%$ for segmentation and detection, respectively. Empirically, we select 512 to be the suitable capacity of the memory storage, not the largest. To this end, the larger capacity can confuse the model in the process of learning when retrieving information in such a large memory bank. Besides, Table IX expresses the effectiveness of the memory loss to the total loss function. As can be inferred, $\beta=1\times 10^{-4}$ gives the best performance evaluated on mAP, AP50, and AP75 among all configurations.
TABLE IX: Ablation study on the $\beta$ ratio of the instance memory loss (Eq. 3) in 1-shot settings. The best, and second best performances are marked in red, and blue, respectively.

### v4

@@ IV-A Overview
As specified in this work, we utilize the proposed CAMO-FS dataset containing images of camouflaged animals in the wild to establish the evaluation of our baseline and proposed improvement. We follow the concept procedure published in FSOD [39, 42, 43]. We employed the MTFA baseline [49] implemented using Detectron2 framework [89]. The backbone is ResNet-101 [90] with Feature Pyramid Network [91]. The models are trained in two stages: base training and novel fine-tuning stage.
In the first stage of the base phase, we train our model with abundant data from $80$ classes with $118K$ images in the train2017 set of the COCO dataset. The training hyper-parameters of the base phase are set according to Detectron2 settings [89].
TABLE V: The number of instances for each camouflaged class in the test set of our CAMO-FS.
Class #ins. Class #ins. Class #ins. Shrimp 10 Crab 37 Dolphin 48 Crocodile 10 Snake 54 Turtle 23 Worm 24 Snail 9 Kangaroo 48 Elephant 12 Giraffe 29 Goat 18 Leopard 39 Monkey 35 Deer 21 Sea_Lion 27 Lion 16 Bat 13 Fox 30 Camel 15 Weasel 32 Rabbit 44 Dog 37 Horse 31 Mouse 8 Hedgehog 10 Rhino 35 Tiger 6 Pig 11 Squirrel 19 Sloth 17 Scorpion 7 Bear 19 Octopus 133 Starfish 71 Seahorse 120 Fish 190 Frog 128 Lizard 180 Cat 82 Moth 56 Penguin 67 Spider 125 Insect 228 Bird 363 Body_Paint. 347 Solider 223 Total 3107
In the second stage of the fine-tuning phase, we evaluate the performance of having $K=\{1,2,3,5\}$ shots per each novel class. Specifically, in the 5-shot setting, we train the novel detector on $47$ camouflaged classes with $197$ images of the CAMO-FS dataset. The training set for other settings is a subset of the 5-shot setting (as presented in Section III-A). The novel phase has a learning rate of $0.00125$ inferred from the MTFA configuration. We set the balance parameters $\alpha$ = $1\times 10^{-1}$ and $\beta$ = $1\times 10^{-2}$ when we train the model with instance triplet loss and instance memory storage, respectively. Other training hyper-parameters of the novel phase are set following TFA [39] settings. Then, the novel models are assessed in a test set including $2,655$ images with $3,107$ instances of $47$ camouflaged classes to obtain results. The number of instances for each class is detailed in Table V. Please visit  [39] or [89] for more details on other parameters of both the training and testing phases. Our models are trained and tested on a single GeForce RTX 2080 Ti GPU. The frame per second (FPS) is approximately 15.
@@
TABLE VI: The improvement of our proposed instance triplet loss (ITL) and instance memory storage (IMS) over the baseline MTFA [39]. The best performances are marked in bold. # denotes the Number of shots.
# Method AP AP50 AP75 APs APm APl AR1 AR10 ARs ARm ARl Instance Segmentation Baseline MTFA 3.66 5.37 4.09 22.42 4.35 2.01 11.30 13.58 25.97 12.96 12.53 MTFA + ITL 4.46 8.21 4.60 21.33 4.13 4.01 12.36 15.04 23.17 9.49 16.67 MTFA + IMS 5.46 9.20 6.17 27.79 6.20 4.01 17.08 19.99 29.41 11.45 20.89 1 MTFA + Both 5.02 8.58 5.38 26.29 5.32 3.93 17.98 21.42 28.13 12.68 22.64 Baseline MTFA 6.21 8.92 7.28 32.64 7.75 3.50 18.88 21.12 35.82 15.49 20.14 MTFA + ITL 5.57 9.45 6.04 25.83 3.01 5.37 15.67 17.33 26.13 7.37 17.50 MTFA + IMS 6.95 10.72 7.60 33.62 5.73 6.44 20.00 22.15 34.25 13.86 20.92 2 MTFA + Both 7.60 11.37 8.26 33.58 5.97 6.49 22.76 24.85 34.02 15.01 24.50 Baseline MTFA 6.16 8.95 6.68 33.74 6.19 5.08 20.25 22.95 36.83 16.31 21.63 MTFA + ITL 6.41 10.67 6.72 30.39 5.17 5.30 20.69 22.98 31.90 15.69 22.53 MTFA + IMS 7.36 11.23 8.49 37.03 6.24 5.64 24.40 27.69 38.44 17.02 26.71 3 MTFA + Both 7.85 11.74 9.05 37.57 5.49 6.69 24.33 27.42 39.05 17.36 25.92 Baseline MTFA 5.95 8.67 6.94 34.71 6.25 4.85 21.29 24.42 36.86 14.51 24.83 MTFA + ITL 8.48 13.43 9.80 36.66 5.75 8.04 23.83 26.66 37.03 11.62 25.91 MTFA + IMS 9.61 14.61 11.73 38.60 5.79 10.40 26.65 30.37 39.21 12.26 30.02 5 MTFA + Both 8.37 13.29 9.45 38.44 5.72 7.60 25.27 29.31 39.33 14.19 28.52 Object Detection Baseline MTFA 2.93 5.86 2.20 20.95 4.18 2.03 9.25 10.84 21.74 11.49 8.77 MTFA + ITL 4.04 8.65 2.98 20.50 4.90 4.22 12.89 15.53 20.73 11.45 17.46 MTFA + IMS 4.50 9.14 3.45 22.88 5.61 3.54 13.14 15.22 23.14 8.78 16.33 1 MTFA + Both 4.31 8.63 3.75 22.15 6.25 3.63 14.67 17.47 22.16 11.62 18.15 Baseline MTFA 5.90 8.87 6.83 33.04 9.74 3.10 17.26 19.25 34.04 15.74 19.61 MTFA + ITL 7.28 11.22 8.25 32.31 10.72 6.83 20.52 22.69 32.34 14.88 23.52 MTFA + IMS 6.95 10.88 7.75 33.93 7.49 6.81 19.84 22.01 34.10 15.04 21.47 2 MTFA + Both 7.60 11.58 8.78 34.13 8.23 6.70 23.87 25.92 34.07 15.13 27.18 Baseline MTFA 5.84 8.98 6.29 34.56 7.78 4.31 19.13 21.83 35.80 15.93 21.09 MTFA + ITL 7.49 11.51 8.23 38.45 8.61 6.38 24.88 27.52 38.55 17.66 27.44 MTFA + IMS 7.55 11.45 8.50 38.07 9.21 5.70 24.20 27.29 38.50 18.10 27.56 3 MTFA + Both 7.94 12.07 9.11 37.97 8.96 6.68 23.77 26.74 38.32 17.64 26.58 Baseline MTFA 5.84 9.13 6.04 35.44 8.17 4.22 19.67 22.96 35.94 14.16 22.58 MTFA + ITL 9.76 14.37 11.12 40.05 8.82 9.89 25.93 29.28 40.05 12.53 30.32 MTFA + IMS 10.36 16.27 11.79 39.32 8.08 11.36 26.34 30.30 39.35 12.37 30.91 5 MTFA + Both 9.39 14.19 10.42 39.37 8.36 9.28 26.16 30.16 39.50 13.98 30.23
@@ IV-B Results and Discussion
State-of-the-art comparison. To prove the effectiveness of our proposed methods, we conducted experiments on our proposed CAMO-FS dataset. We tested with $K=\{1,2,3,5\}$ shots, respectively. Since several recent work have not published their source code [82, 81], we adopted the typical models addressing both detection and instance segmentation tasks to compare with our proposed methods. Table IV presents the evaluation of the performance of our methods of instance triplet loss and memory storage over our baseline MTFA [49], the model of Mask R-CNN [66] with sigmoid classifier, and the state-of-the-art method iFS-RCNN [79] in the approach of few-shot instance segmentation. We reported experiments on those models and chose the common COCO-80 ResNet-101 as their base model to apply our proposed methods. The details of this decision are declared in the ablation section. In terms of instance segmentation, we improved over MTFA [49], Mask RCNN [66], and iFS-RCNN [79] by getting average AP values of $6.23\%,8.04\%,7.13\%$, respectively thanks to instance triplet loss, and $7.35\%,7.96\%,6.25\%$, respectively thanks to instance memory storage. Regarding object detection, our FS-CDIS got average amounts of $7.14\%,7.34\%,6.96\%$, respectively with instance triplet loss and $7.34\%,7.45\%,5.88\%$, respectively with instance memory storage. The detailed performance of our methods is in Table IV. Despite the limited results, we defeated the very early models on detection and instance segmentation tasks on camouflaged images.
@@
Proposed modules evaluation. In Table VI, we also present the results of the baseline MTFA [49] with its original default configuration along with our proposed improvements. On top of the baseline MTFA [49], we establish fine-tuning configuration on this model by training all heads of classification, box regression, and mask prediction on few-shot novel data. The reported results prove the performance of the proposed instance triplet loss, instance memory storage, and the combination of both loss functions.
@@
TABLE VII: Ablation study on the base model with 1-shot results. The best performances are marked in bold. “Triplet” stands for Instance Triplet Loss and “Memory” stands for Instance Memory Storage.
@@
In general, our approaches achieve outstanding results in comparison with the baseline. Our improvements surpass MTFA by a remarkable margin. These results manifest the efficiency of our methods in the context of few-shot camouflaged instance segmentation. Both loss functions enhance the discrimination between foreground and background features which strongly supports the model to segment pixels that belong to the camouflaged animals. Regarding the memory storage and the triplet loss function, the results of the memory loss function are higher than those of the triplet loss function by about $1\%$. We realize that storing representatives for each class is a crucial element in few-shot learning. This technique not only expands the variants during training but also increases the consistency per class, so thereby model can segment difficult objects better. In these ways, we also improve the corresponding results in camouflage object detection.
@@
In Table VI, our improvements help the model segment animals in various sizes. Specifically, all three metrics including APs, APm, and APl improve in comparison with the baseline model, which demonstrates that our model well segments small, medium, and large animals. This situation also happens in the detection task. When data is very scarce as in a 1-shot or 2-shot setting, the instance triplet loss function has comparative results with the instance memory storage function. However, in the context of 3-shot or 5-shot settings, the instance memory storage demonstrates outstanding efficiency thanks to storing and updating the memory via iterations to create discriminative features on a global level. Actually, our proposed instance triplet loss is designed to differentiate the features among the foreground and background of a single camouflaged instance. Meanwhile, the instance memory storage aims to store features of multiple instances of the same category to enhance the general features. Thus, when combining the two approaches at the same time to train our FS-CDIS, the performance of the model fluctuates but still follows a trend. It can be seen from the reported numbers that with $K=\{1,2,3\}$ shots, the results of the combined loss seem to dominate the results of each separate component. However, notably around $K=5$ shots, the AP, AP50, and AP75 of the instance memory storage yield better performance due to the information increase by more shots.
In terms of quantitative comparison, Figure 8 illustrates the qualitative comparison among the results of 5-shot settings of the baseline MTFA [49] and our proposed methods of Instance Triplet Loss and Instance Memory Storage. We chose to visualize the images with a confidence threshold of $0.5$, which released a huge number of predictions with low confidence from the models. The four final rows indicate exemplary cases that either FS-CDIS-ITL or FS-CDIS-IMS can figure out camouflaged instances compared to the baseline. In these cases, our methods seem to be better the baseline although there are some imperfect cases where the prediction mask in the sixth row overlays irrelevant parts of the object (over-segmentation) or the results in the fourth, fifth, and last rows only capture some main parts of objects (under-segmentation). We conjecture the reasons for that is the image contains background clutter and the extremely vague boundary between foreground and background. To alleviate the phenomenon, it is feasible to further apply the post-processing methods on the output segmentation masks.
@@
TABLE VIII: Ablation study on the margin and the $\alpha$ ratio of the instance triplet loss in 1-shot settings. The best performances are marked in bold.
@@
Base model ablation study. We also conduct ablation experiments on different backbone base models of the COCO settings including general and few-shot concepts. To be detailed, we report the performance of our proposed method of instance triplet loss and instance memory storage over four different backbones. The considered backbones are ResNet-50 and ResNet-101 [90]. The two base datasets are MS-COCO with 80 classes and 60 classes, respectively. Thus, it led to the combination of four different base models (i.e. COCO-80/60 R-101, COCO-80/60 R-50). As can be seen from Table VII, the performance of applying COCO-80 R-101 base weight yields better results among others evaluated on AP, AP@50, and AP@75 in both segmentation and detection tasks. In both cases of our two proposed improvements, the ablation results demonstrate our selection of COCO-80 R-101 is the best among the tested backbones of the base phase. For the segmentation task, we achieve $4.46$ and $5.46$ of AP reported for triplet loss and memory storage, respectively. For the detection task, we reach $4.04$ and $4.50$ also of AP, respectively. In summary, the chosen backbone of the base weight presents a higher performance of around $1\%$ to $2\%$ of evaluated on common metrics as in the table. To be explained, the base from COCO-80 contains more semantic concepts in comparison with the COCO-60 base, which leads to the higher performance reported. Note that all the results in this ablation section are reported for the 1-shot setting.
Ablation on instance triplet loss component. In terms of the instance triplet loss described in Eq. 1, we establish ablation experiments to evaluate the performance of the model with different configurations of margin and $\alpha$ value when the instance memory storage component is disabled (i.e. $\beta=0$). To this end, we set up the margin varying from 00 to $1$, with a step of $0.25$. For the $\alpha$ ratio of the loss function (Eq. 3), we check out $\alpha$ = {$1,1\times 10^{-1},1\times 10^{-2}$}. To be enhanced, the margin value indicates how distinguished foreground and background features are. Meanwhile, the $\alpha$ controls the effect of the instance triplet loss on the total loss of the framework. Table VIII presents the evaluation of both detection and segmentation issues in 1-shot manner. As can be inferred from the table, the effect of $\alpha$ decides which margin should be selected for the triplet loss. With $\alpha=1$ meaning we keep the original ratio of the loss, the segmentation result in 1-shot setting yields the highest performance of $5.16\%$ mAP with a $0.75$ margin value. Meanwhile, the detection result gets the highest performance of $4.36\%$ with $\alpha=1\times 10^{-1}$ and zero margin. This table offers a better understanding of the impact of $\alpha$ and the margin over the total performance.
TABLE IX: Ablation study on the capacity of the instance memory storage in 1-shot settings. The best performances are marked in bold.
@@
Ablation on instance memory storage component. As for the instance memory storage, as introduced in Eq. 2, and Eq. 3, there are several parameters that need analyzing, listed as the amount of capacity in the memory storage and the $\beta$ ratio controlling the effect of the memory storage loss in the total loss. Table IX, and Table X present the ablation experimental results of those issues when the instance triplet loss function is disabled (i.e. $\alpha=0$), respectively. In terms of the capacity of the memory storage, we establish experiments on a range of memory capacity of $2^{i}$ where $i=\{5,6,7,8,9,10\}$. The reported results figure out that the performance on both segmentation and detection tasks increases with a larger capacity of memory storage. To be detailed, with a capacity of $512$, the mAP metric achieves the highest value among configurations, i.e. $4.76\%$ and $4.48\%$ for segmentation and detection, respectively. Empirically, we select $512$ to be the suitable capacity of the memory storage, not the largest. To this end, the larger capacity can confuse the model in the process of learning when retrieving information in such a large memory bank. Besides, Table X expresses the effectiveness of the memory loss to the total loss function. As can be inferred, $\beta=1\times 10^{-4}$ gives the best performance evaluated on mAP, AP50, and AP75 among all configurations.
TABLE X: Ablation study on the $\beta$ ratio of the instance memory loss (Eq. 3) in 1-shot settings. The best performances are marked in bold.